# Advancing Image Super-resolution Techniques in Remote Sensing: A Comprehensive Survey

### v2

@@ 4.1.4 Probability-based Methods
Probability-based methods mainly utilizes statistical modeling and probabilistic inference to perform image reconstruction. Specifically, these methods inherently incorporate prior knowledge of the imaging scene, enabling unsupervised reconstruction without relying on paired high-low-resolution training datasets. Generally, probabilistic RSISR methods can be systematically categorized into four paradigms: Bayesian inference, maximum a posteriori (MAP) estimation, maximum likelihood (ML) estimation, and Markov random field (MRF) modeling.
@@
Bayesian-based Methods. Generally speaking, Bayesian-based methods transform the ill-posed problem of super-resolution reconstruction into a well-posed inversion model by combining image prior probability density with physical constraints. In the past two decades, Bayesian RSISR has been widely studied, and many Bayesian RSISR methods have been proposed. To the best of our knowledge, Tipping and Bishop [288] first proposed a Bayesian super-resolution method based on marginalization in 2002. Specifically, they modeled high-resolution images by injecting Gaussian process priors to achieve joint estimation of point spread function and registration parameters. Besides, Molina et al. [289] integrated the sensor characteristics into the Bayesian framework and combined the observation process of multispectral and panchromatic images to effectively improve the spatial resolution of Landsat ETM+ while maintaining the spectral fidelity. Subsequently, Molina et al. [290, 291, 292] further constructed a hierarchical Bayesian model with parameter priors and used variational inference methods to achieve simultaneous estimation of HR multispectral images and model parameters. In addition, Babacan et al. [127] proposed a joint super-resolution reconstruction method based on variational Bayesian, which simultaneously estimates high-resolution images and motion parameters through probabilistic modeling to achieve adaptive robust reconstruction. For pansharpening, Wang et al. [293] proposed a Bayesian fusion model based on the geometric-spectral-spatial consistency hypothesis, which uses the alternating direction multiplier method (ADMM) to solve and verify the generalization ability of hyperspectral images. In order to solve the problem of parameter sensitivity, Armannsson et al. [294] innovatively introduced Bayesian optimization into Sentinel-2 super-resolution, and significantly improved the robustness of methods such as ATPRK and S2Sharp through a small-scale parameter optimization mechanism. Li et al. [295] proposed a multi-prior Bayesian model based on the fusion of TV and Laplace priors, which significantly improved the accuracy of target contour reconstruction in sea and land scenes. Furthermore, Tan et al. [296] established a two-dimensional spatial structure prior based on Markov random field (MRF), and characterized the spatial correlation of the scene through an n-order neighborhood system, which significantly improved the suppression of artifacts compared with previous methods. Besides, Ye et al. [81] proposed a Bayesian hyperspectral super-resolution method based on texture decomposition, which effectively suppressed spectral variation through Gaussian process spectral prior. In recent work, Shen et al. [297] and Guo et al. [113] achieved breakthroughs in RSISR of Bayesian forward-looking radar imaging in low signal-to-noise ratio and maneuvering scenarios by using gamma-normal conjugate model and structured Student’s t prior, respectively.
@@
MAP-based Methods. RSISR methods based on maximum a posteriori probability (MAP) have also been widely studied. For example, the early work of Chantas et al. [298] proposed a MAP framework with local adaptive edge-preserving priors, which achieved joint restoration and registration under degraded observations through a two-step reconstruction strategy. Subsequently, some studies began to optimize from the perspective of prior models. Wang et al. [299] proposed a MAP method based on high-frequency fusion of panchromatic images for RSISR, which introduced spectral fidelity prior constraints to iteratively optimize the framework. In addition, Belekos et al. [300] designed a multi-channel non-stationary Gaussian prior model to enhance the collaborative reconstruction capability of multispectral data. Li et al. [301] first introduced the Hidden Markov tree (HMT) into the MAP framework and used the multi-scale dependence of wavelet coefficients to improve the reconstruction performance. Improved methods for specific application scenarios have also emerged. For example, Guan et al. [302] proposed a MAP radar angle super-resolution method based on antenna pattern and target scattering convolution prior modeling. Zhang et al. [63] proposed a Rayleigh sparse MAP (RSMAP) algorithm for sea surface target imaging, which combines the distribution characteristics of sea clutter with the sparse prior of the target to improve the angular resolution. In terms of noise robustness, Vrigkas et al. designed a global M-estimation framework to achieve parameter adaptive adjustment. Besides, the regional adaptive total variation (RSATV) model proposed by Yuan et al. [303] for RSISR utilized spatial information classification to suppress artifacts. On the other hand, Li et al. [304] first achieved RSISR on short-time sequence images of the GF-4 geosynchronous satellite, verifying the feasibility of synchronous orbit observation data. In addition, some studies have constructed the MAP-MRF framework to perform hyperspectral super-resolution technology. Irmak et al. [305] achieved multi-frame reconstruction optimization through abundance map decomposition and endmember extraction. Further, the extended study [56] combined virtual dimension endmember determination with texture constrained secondary optimization, and its performance exceeded the mainstream method while maintaining spectral consistency, forming a complete technical chain for multi-source to single-source reconstruction.
@@
MRF-based Methods. The Markov Random Field (MRF) based method mainly models the spatial dependencies between pixels and their neighborhoods, which performs effective integration of global spatial context information through the energy function of the local neighborhood. MRF-based RSISR has been studied for nearly two decades, and many MRF-based RSISR methods have been proposed. For example, Li et al. [309] proposed a spatial-temporal MRF super-resolution model (STMRF_SRM), which integrates historical medium-resolution images with current coarse-resolution data to effectively enhance the spatiotemporal consistency of forest mapping. Aghighi et al. [310] further proposed a full-space adaptive MRF-SRM, which dynamically balances spectral and spatial weights based on endmember analysis and local energy matrix to achieve automatic parameter calibration. In the latest progress, Welikanna et al. [311] fused the fuzzy C-means and spectral angle quantity to optimize the MRF posterior energy function, which reduces the influence of point spread effect.
@@ 4.2 Deep learning-based RSISR Method
The majority of deep learning-based RSISR methods are supervised, which need a large amount of labeled training data. However, the ground truth is normally not available in RSISR, thus almost all supervised methods utilize synthetic degraded simulation datasets that are actually not realistic. Therefore, it is desirable to have unsupervised RSISR methods. In this section, we investigate the unsupervised deep learning RSISR methods, mainly involving self-supervised learning, contrastive learning, generative-based and zero-shot learning.
@@ 4.2.1 Self-supervised Learning
Self-supervised RSISR methods are usually inspired by internal degradation modeling or cross-modal data association to generate pseudo-supervisory signals, thereby avoiding the reliance on external HR training data. In recent years, self-supervised RSISR methods have emerged. In 2018, Haut et al. [54] proposed the first unsupervised convolutional generation model, which learned the LR-HR feature relationship through stacked convolution and downsampling layers, and employed LR reconstruction loss to constrain the generation quality. In addition, Choi and Kim [312] proposed a new degradation model to generate LR samples, and combined it with morphological transformation to enhance HR images, thereby achieving super-resolution of KOMPSAT-3 images. Sheikholeslami et al. [68] proposed an efficient unsupervised super-resolution (EUSR) model for RSISR, which uses dense skip connections and bottleneck compression modules to reduce the amount of computation while maintaining reconstruction performance. For cross-modal data, Chen et al. [74] proposed a self-supervised spectral-spatial residual network (SSRN), which generates HR hyperspectral images through the mapping relationship between LR multispectral and hyperspectral images without HR ground truth supervision. In terms of multi-frame super-resolution, Nguyen et al. [313] proposed an frame-to-frame self-supervised framework that fuses satellite image sequence features and improves the resolution of SkySat satellite images. In the latest progress, Hong et al. [314] proposed a decoupled-and-coupled network (DC-Net) for RSISR, which effectively alleviated the distribution difference of hyperspectral and multispectral data through the decoupled-coupled structure and self-supervised constraints. Qian et al. [39] proposed a RSISR method based on deep image prior. Specifically, they used convolutional neural networks and 3D separable convolutions to perform super-resolution restoration on the coarse resolution band of Sentinel-2 satellite images without the need for additional training data, significantly improving the super-resolution performance. An interesting work is a self-supervised degradation guidance method proposed by Xiao et al. [89], which effectively adapts to various unknown degradations through contrastive learning and bidirectional feature modulation network, significantly improving the RSISR performance.
@@ 4.2.2 Contrastive Learning
Traditional methods rely on bicubic downsampling to synthesize LR images for training, which is difficult to reflect the complex degradation patterns of real scenes. Contrastive learning, as a self-supervised learning method, unsupervisedly compares the degradation features and HR detail differences between samples to mine potential degradation representations, significantly improving the model’s adaptability to unknown degradation and the physical consistency of super-resolution reconstruction. In recent years, some RSISR methods based on contrastive learning have been proposed. Mishra and Hadar [88] proposed a semi-supervised contrastive learning framework CLSR, which significantly reduced the reliance on large-scale diverse training samples by comparing the potential differences between real degradation and ideal bicubic degradation. Further, the extended research [315] proposed a two-stage contrast framework for RSISR. In the first stage, ”artificial style images” with HR textures are generated through contrast training, and then neural style transfer is combined to achieve unsupervised super-resolution. An interesting contrastive learning-based RSISR method is the region-aware network (RAN) proposed by Liu et al. [87]. In RAN, they integrated the degradation prior extracted by contrastive learning with the regional structure analysis module, and captured cross-block self-similarity through graph neural networks and attention mechanisms, effectively improving the fidelity of texture reconstruction. In the latest progress, the EDRLN network proposed by Wang et al. [106] uses mutual affine convolution to replace the traditional convolution layer and integrates a lightweight pixel attention mechanism in the contrastive learning framework, which reduces the number of model parameters while maintaining excellent performance in multiple degradation scenarios.
@@ 4.2.3 Generative-based Methods
Unsupervised generative methods establish data distribution mappings rather than relying on paired samples to perform RSISR. In recent years, some studies have used GAN or cyclic structures to learn the joint distribution of degradation and reconstruction, and achieve resolution enhancement through implicit image priors. For example, Wang et al. [316] proposed a Cycle-CNN based on a cyclic structure. They used two generative CNNs to model the downsampling and super-resolution processes respectively, and verified the effectiveness of unpaired training on GaoFen-2 satellite data. In 2021, the extended research [72] proposed a new Cycle-CNN that integrates the bidirectional mapping of the downsampling network and the SR network, achieving improved robustness on the UC Merced dataset, and further demonstrated the ability of the cyclic generative structure to suppress noise and blur. In addition, Wang et al. [73] proposed an enhanced image prior (EIP) method for RSISR, which encodes the reference image into the latent space based on GAN, and combines the cyclic update strategy to transfer texture features, achieving advanced performance. Similarly, the MIP method [71] constructs a transfer image prior and uses an implicit noise update mechanism to achieve feature transfer for RSISR. Zhang et al. [317] proposed an unsupervised RSISR method based on GAN, which achieved competitive performance by reconstructing SR images through generator and then downsampling to train the discriminator. A recent work is TransCycleGAN proposed by Zhai et al. [107], which integrates the Transformer module into the CycleGAN framework and uses transposed self-attention to capture global context, effectively removing degenerate features in a pseudo-supervisory manner.

### v3

@@ 4.1.4 Probability-based Methods
Probability-based methods mainly utilize statistical modeling and probabilistic inference to perform image reconstruction. Specifically, these methods inherently incorporate prior knowledge of the imaging scene, enabling unsupervised reconstruction without relying on paired high-low-resolution training datasets. Generally, probabilistic RSISR methods can be systematically categorized into four paradigms: Bayesian inference, maximum a posteriori (MAP) estimation, maximum likelihood (ML) estimation, and Markov random field (MRF) modeling.
@@
Bayesian-based Methods. Generally speaking, Bayesian-based methods transform the ill-posed problem of super-resolution reconstruction into a well-posed inversion model by combining image prior probability density with physical constraints. In the past two decades, Bayesian RSISR has been widely studied, and many Bayesian RSISR methods have been proposed. To the best of our knowledge, Tipping and Bishop [291] first proposed a Bayesian super-resolution method based on marginalization in 2002. Specifically, they modeled high-resolution images by injecting Gaussian process priors to achieve joint estimation of point spread function and registration parameters. Besides, Molina et al. [292] integrated the sensor characteristics into the Bayesian framework and combined the observation process of multispectral and panchromatic images to effectively improve the spatial resolution of Landsat ETM+ while maintaining the spectral fidelity. Subsequently, Molina et al. [293, 294, 295] further constructed a hierarchical Bayesian model with parameter priors and used variational inference methods to achieve simultaneous estimation of HR multispectral images and model parameters. In addition, Babacan et al. [129] proposed a joint super-resolution reconstruction method based on variational Bayesian, which simultaneously estimates high-resolution images and motion parameters through probabilistic modeling to achieve adaptive robust reconstruction. For pansharpening, Wang et al. [296] proposed a Bayesian fusion model based on the geometric-spectral-spatial consistency hypothesis, which uses the alternating direction multiplier method (ADMM) to solve and verify the generalization ability of hyperspectral images. In order to solve the problem of parameter sensitivity, Armannsson et al. [297] innovatively introduced Bayesian optimization into Sentinel-2 super-resolution, and significantly improved the robustness of methods such as ATPRK and S2Sharp through a small-scale parameter optimization mechanism. Li et al. [298] proposed a multi-prior Bayesian model based on the fusion of TV and Laplace priors, which significantly improved the accuracy of target contour reconstruction in sea and land scenes. Furthermore, Tan et al. [299] established a two-dimensional spatial structure prior based on Markov random field (MRF), and characterized the spatial correlation of the scene through an n-order neighborhood system, which significantly improved the suppression of artifacts compared with previous methods. Besides, Ye et al. [82] proposed a Bayesian hyperspectral super-resolution method based on texture decomposition, which effectively suppressed spectral variation through Gaussian process spectral prior. In recent work, Shen et al. [300] and Guo et al. [114] achieved breakthroughs in RSISR of Bayesian forward-looking radar imaging in low signal-to-noise ratio and maneuvering scenarios by using a gamma-normal conjugate model and structured Student’s t prior, respectively.
@@
MAP-based Methods. RSISR methods based on maximum a posteriori probability (MAP) have also been widely studied. For example, the early work of Chantas et al. [301] proposed a MAP framework with local adaptive edge-preserving priors, which achieved joint restoration and registration under degraded observations through a two-step reconstruction strategy. Subsequently, some studies began to optimize from the perspective of prior models. Wang et al. [302] proposed a MAP method based on high-frequency fusion of panchromatic images for RSISR, which introduced spectral fidelity prior constraints to iteratively optimize the framework. In addition, Belekos et al. [303] designed a multi-channel non-stationary Gaussian prior model to enhance the collaborative reconstruction capability of multispectral data. Li et al. [304] first introduced the Hidden Markov tree (HMT) into the MAP framework and used the multi-scale dependence of wavelet coefficients to improve the reconstruction performance. Improved methods for specific application scenarios have also emerged. For example, Guan et al. [305] proposed a MAP radar angle super-resolution method based on antenna pattern and target scattering convolution prior modeling. Zhang et al. [64] proposed a Rayleigh sparse MAP (RSMAP) algorithm for sea surface target imaging, which combines the distribution characteristics of sea clutter with the sparse prior of the target to improve the angular resolution. In terms of noise robustness, Vrigkas et al. designed a global M-estimation framework to achieve parameter adaptive adjustment. Besides, the regional adaptive total variation (RSATV) model proposed by Yuan et al. [306] for RSISR utilized spatial information classification to suppress artifacts. On the other hand, Li et al. [307] first achieved RSISR on short-time sequence images of the GF-4 geosynchronous satellite, verifying the feasibility of synchronous orbit observation data. In addition, some studies have constructed the MAP-MRF framework to perform hyperspectral super-resolution technology. Irmak et al. [308] achieved multi-frame reconstruction optimization through abundance map decomposition and endmember extraction. Further, the extended study [57] combined virtual dimension endmember determination with texture-constrained secondary optimization, and its performance exceeded the mainstream method while maintaining spectral consistency, forming a complete technical chain for multi-source to single-source reconstruction.
@@
MRF-based Methods. The Markov Random Field (MRF) based method mainly models the spatial dependencies between pixels and their neighborhoods, which performs effective integration of global spatial context information through the energy function of the local neighborhood. MRF-based RSISR has been studied for nearly two decades, and many MRF-based RSISR methods have been proposed. For example, Li et al. [312] proposed a spatial-temporal MRF super-resolution model (STMRF_SRM), which integrates historical medium-resolution images with current coarse-resolution data to effectively enhance the spatiotemporal consistency of forest mapping. Aghighi et al. [313] further proposed a full-space adaptive MRF-SRM, which dynamically balances spectral and spatial weights based on endmember analysis and local energy matrix to achieve automatic parameter calibration. In the latest progress, Welikanna et al. [314] fused the fuzzy C-means and spectral angle quantity to optimize the MRF posterior energy function, which reduces the influence of the point spread effect.
@@ 4.2 Deep learning-based RSISR Method
The majority of deep learning-based RSISR methods are supervised, which require a large amount of labeled training data. However, the ground truth is normally not available in RSISR, thus, almost all supervised methods utilize synthetic degraded simulation datasets that are actually not realistic. Therefore, it is desirable to have unsupervised RSISR methods. In this section, we investigate the unsupervised deep learning RSISR methods, mainly involving self-supervised learning, contrastive learning, generative-based, and zero-shot learning.
@@ 4.2.1 Self-supervised Learning
Self-supervised RSISR methods are usually inspired by internal degradation modeling or cross-modal data association to generate pseudo-supervisory signals, thereby avoiding the reliance on external HR training data. In recent years, self-supervised RSISR methods have emerged. In 2018, Haut et al. [55] proposed the first unsupervised convolutional generation model, which learned the LR-HR feature relationship through stacked convolution and downsampling layers, and employed LR reconstruction loss to constrain the generation quality. In addition, Choi and Kim [315] proposed a new degradation model to generate LR samples, and combined it with morphological transformation to enhance HR images, thereby achieving super-resolution of KOMPSAT-3 images. Sheikholeslami et al. [69] proposed an efficient unsupervised super-resolution (EUSR) model for RSISR, which uses dense skip connections and bottleneck compression modules to reduce the amount of computation while maintaining reconstruction performance. For cross-modal data, Chen et al. [75] proposed a self-supervised spectral-spatial residual network (SSRN), which generates HR hyperspectral images through the mapping relationship between LR multispectral and hyperspectral images without HR ground truth supervision. In terms of multi-frame super-resolution, Nguyen et al. [316] proposed a frame-to-frame self-supervised framework that fuses satellite image sequence features and improves the resolution of SkySat satellite images. In the latest progress, Hong et al. [317] proposed a decoupled-and-coupled network (DC-Net) for RSISR, which effectively alleviated the distribution difference of hyperspectral and multispectral data through the decoupled-coupled structure and self-supervised constraints. Qian et al. [40] proposed an RSISR method based ona deep image prior. Specifically, they used convolutional neural networks and 3D separable convolutions to perform super-resolution restoration on the coarse resolution band of Sentinel-2 satellite images without the need for additional training data, significantly improving the super-resolution performance. An interesting work is a self-supervised degradation guidance method proposed by Xiao et al. [90], which effectively adapts to various unknown degradations through contrastive learning and a bidirectional feature modulation network, significantly improving the RSISR performance.
@@ 4.2.2 Contrastive Learning
Traditional methods rely on bicubic downsampling to synthesize LR images for training, which cannot reflect the complex degradation patterns of real scenes. Contrastive learning, as a self-supervised learning method, unsupervisedly compares the degradation features and HR detail differences between samples to mine potential degradation representations, significantly improving the model’s adaptability to unknown degradation and the physical consistency of super-resolution reconstruction. In recent years, some RSISR methods based on contrastive learning have been proposed. Mishra and Hadar [89] proposed a semi-supervised contrastive learning framework termed CLSR, which significantly reduced the reliance on large-scale diverse training samples by comparing the potential differences between real degradation and ideal bicubic degradation. Further, the extended research [318] proposed a two-stage contrast framework for RSISR. In the first stage, ”artificial style images” with HR textures are generated through contrast training, and then neural style transfer is combined to achieve unsupervised super-resolution. An interesting contrastive learning-based RSISR method is the region-aware network (RAN) proposed by Liu et al. [88]. In RAN, they integrated the degradation prior extracted by contrastive learning with the regional structure analysis module, and captured cross-block self-similarity through graph neural networks and attention mechanisms, effectively improving the fidelity of texture reconstruction. In the latest progress, the EDRLN network proposed by Wang et al. [107] uses mutual affine convolution to replace the traditional convolution layer and integrates a lightweight pixel attention mechanism in the contrastive learning framework, which reduces the number of model parameters while maintaining excellent performance in multiple degradation scenarios.
@@ 4.2.3 Generative-based Methods
Unsupervised generative methods establish data distribution mappings rather than relying on paired samples to perform RSISR. In recent years, some studies have used GAN or cyclic structures to learn the joint distribution of degradation and reconstruction, and achieve resolution enhancement through implicit image priors. For example, Wang et al. [319] proposed a Cycle-CNN based on a cyclic structure. They used two generative CNNs to model the downsampling and super-resolution processes, respectively, and verified the effectiveness of unpaired training on GaoFen-2 satellite data. In 2021, the extended research [73] proposed a new Cycle-CNN that integrates the bidirectional mapping of the downsampling network and the SR network, achieving improved robustness on the UC Merced dataset, and further demonstrated the ability of the cyclic generative structure to suppress noise and blur. In addition, Wang et al. [74] proposed an enhanced image prior (EIP) method for RSISR, which encodes the reference image into the latent space based on GAN, and combines the cyclic update strategy to transfer texture features, achieving advanced performance. Similarly, the MIP method [72] constructs a transfer image prior and uses an implicit noise update mechanism to achieve feature transfer for RSISR. Zhang et al. [320] proposed an unsupervised RSISR method based on GAN, which achieved competitive performance by reconstructing SR images through a generator and then downsampling to train the discriminator. A recent work is TransCycleGAN proposed by Zhai et al. [108], which integrates the Transformer module into the CycleGAN framework and uses transposed self-attention to capture global context, effectively removing degenerate features in a pseudo-supervisory manner.